### v1

# neurips_2025.bib

Abstract
Many architectures for neural PDE surrogates have been proposed in recent years, largely based on neural networks or operator learning. In this work, we derive and propose a new architecture, the Neural Functional, which learns function to scalar mappings. Its implementation leverages insights from operator learning and neural fields, and we show the ability of neural functionals to implicitly learn functional derivatives. For the first time, this allows for an extension of Hamiltonian mechanics to neural PDE surrogates by learning the Hamiltonian functional and optimizing its functional derivatives. We demonstrate that the Hamiltonian Neural Functional can be an effective surrogate model through improved stability and conserving energy-like quantities on 1D and 2D PDEs. Beyond PDEs, functionals are prevalent in physics; functional approximation and learning with its gradients may find other uses, such as in molecular dynamics or design optimization.
1 Introduction
Physics is remarkably complex; a set of governing laws can produce endlessly diverse and interesting phenomena. In an effort to unify governing laws, Hamiltonian mechanics describes physical phenomena under the Principle of Least Action, whereby the trajectory of a system evolves because it has locally the least action. Remarkably, the minimization of action can derive nearly all physical phenomena, from kinematics to quantum mechanics, suggesting an underlying, fundamental property of nature. Rather than describing physics through forces and masses, such as in Newtonian mechanics, using energy and symmetries allows for an elegant and unifying perspective of physical systems.
Despite the long history of Hamiltonian mechanics, most current numerical and neural PDE solvers operate within a Newtonian framework. This has been very successful; forces and masses are easy to intuitively understand, and both numerical and neural solvers have found important, practical uses. However, motivated by natural conservation laws and symmetries that arise from Hamiltonian mechanics, we seek to investigate neural PDE surrogates that operate in this framework. We find that the Hamiltonian framework brings inductive biases that improve model performance, especially when conserving energy-like quantities.
Extending Hamiltonian mechanics to neural PDE surrogates requires both theoretical and implementation insights. In particular, most PDEs are used to describe continuum systems (fluids, waves, elastic bodies, etc.); as such, infinite-dimensional Hamiltonian mechanics are needed. Within this setup, the Hamiltonian is defined as a functional that maps from input fields to a scalar, often interpreted as the energy of the system. The variational or functional derivative of the Hamiltonian then contains dynamical information that evolves the system. This introduces new modeling challenges; in particular, no current neural architecture can theoretically model this infinite-dimensional to scalar mapping, let alone have meaningful functional derivatives.
In this paper, we introduce the Neural Functional, an architecture that can provably represent functionals in the linear case and implicitly learn functional derivatives. Through automatic differentiation, its functional derivative can obtained and optimized in a Hamiltonian framework to predict future PDE states. We evaluate the Hamiltonian Neural Functional (HNF) on three PDE systems and find both high accuracy and generalization capabilities, which we hypothesize arise from inductive biases to conserve energy. Beyond predicting PDEs, functionals arise in many physical systems, opening interesting opportunities to apply and improve neural functionals for other meaningful uses. Code and datasets for this work are released at https://github.com/anthonyzhou-1/hamiltonian_pdes.
2 Background
Hamiltonian Mechanics
Hamiltonian mechanics is usually introduced in the discrete setting, where there are a set of $n$ bodies, each with a position and momentum. Therefore, the state of the system can be described by $2n$ quantities, usually assembled into a position and momentum vector $\mathbf{q},\mathbf{p}\in\mathbb{R}^{n}$. In the discrete case, the Hamiltonian $\{\mathcal{H}(\mathbf{q},\mathbf{p}):\mathbb{R}^{2n}\rightarrow\mathbb{R}\}$ is a function that takes two vectors and returns a scalar $\mathcal{H}$, usually interpreted as the energy or another conserved quantity. Evolving the system in time is done through Hamilton’s equations, often expressed in terms of the symplectic matrix $J$ and state vector $\mathbf{u}$:
$\underbrace{\frac{\text{d}\mathbf{q}}{\text{d}t}=\frac{\partial\mathcal{H}}{\partial\mathbf{p}},\quad\frac{\text{d}\mathbf{p}}{\text{d}t}=-\frac{\partial\mathcal{H}}{\partial\mathbf{q}}}_{\text{Hamilton's Equations}}\qquad\underbrace{\frac{\text{d}\mathbf{u}}{\text{d}t}=\begin{bmatrix}\text{d}\mathbf{q}/\text{d}t\\ \text{d}\mathbf{p}/\text{d}t\end{bmatrix}=\begin{bmatrix}0&\mathbb{I}\\ -\mathbb{I}&0\end{bmatrix}\begin{bmatrix}\partial\mathcal{H}/\partial\mathbf{q}\\ \partial\mathcal{H}/\partial\mathbf{p}\end{bmatrix}=J\nabla_{\mathbf{u}}\mathcal{H}}_{\text{Symplectic Form}}$ (1)
Derivatives can be evaluated from vector calculus identities; for clarity we consider the state vector $\mathbf{u}$ with $2n$ components and write $\frac{\text{d}\mathbf{u}}{\text{d}t}=[\frac{\text{d}u_{1}}{\text{d}t},\ldots,\frac{\text{d}u_{2n}}{\text{d}t}]$, and $\nabla_{\mathbf{u}}\mathcal{H}=[\frac{\partial\mathcal{H}}{\partial u_{1}},\ldots,\frac{\partial\mathcal{H}}{\partial u_{2n}}]$.
Infinite-Dimensional Systems
Many PDEs describe continuum systems, such as waves, fluids, and elastic bodies, where the positions and momenta of discrete bodies are not well defined. Despite this change, continuum systems can still conserve energy and be viewed in a Hamiltonian framework. Firstly, finite-dimensional vectors $\mathbf{q},\mathbf{p}\in\mathbb{R}^{n}$ become generalized to functions $u(x,t)$ defined on continuously varying coordinates $x\in\Omega$ and at a time $t$. The function $u\in\mathcal{F}(\Omega)$ describes the state of the system by assigning a scalar or vector for each coordinate (such as velocity, pressure, etc.). Secondly, the Hamiltonian is generalized from a function $\{\mathcal{H}(\mathbf{q},\mathbf{p}):\mathbb{R}^{2n}\rightarrow\mathbb{R}\}$ to a functional $\{\mathcal{H}[u]:\mathcal{F}(\Omega)\rightarrow\mathbb{R}\}$ that returns a scalar given an input function. Lastly, the symplectic matrix $J$ generalizes to a linear operator $\mathcal{J}$, and the gradient $\nabla_{\mathbf{u}}$ becomes the variational or functional derivative $\frac{\delta\mathcal{H}}{\delta u}\in\mathcal{F}(\Omega)$. This leads to Hamilton’s equations for infinite-dimensional systems, where $\epsilon\in\mathcal{F}(\Omega)$ is an arbitrary test function:
$\underbrace{\frac{\partial u}{\partial t}=\mathcal{J}\frac{\delta\mathcal{H}}{\delta u}}_{\text{Hamilton's Equations}}\qquad\underbrace{\int_{\Omega}\frac{\delta\mathcal{H}}{\delta u}(x)\epsilon(x)\text{d}x=\lim_{h\rightarrow 0}\frac{\mathcal{H}[u+h\epsilon]-\mathcal{H}[u]}{h}}_{\text{Functional Derivative}}$ (2)
[…]
3 Neural Functionals
Theory
The development and underlying theory of neural functionals is a result of the Riesz representation theorem, a fundamental result in functional analysis [kreyszig1991introductory]. We give an informal derivation of neural functionals to provide intuition for the architecture; a full proof is in Appendix C. We start by observing that a functional can be equivalent to the inner product over suitable functions:
Theorem 3.1 (Riesz representation theorem).
Let $H$ be a Hilbert space whose inner product $\left\langle x,y\right\rangle$ is defined. For every continuous linear functional $\varphi\in H^{*}$ there exists a unique function $f_{\varphi}\in H$, called the Riesz representation of $\varphi$, such that:
$\varphi[x]=\left\langle x,f_{\varphi}\right\rangle\quad{\text{ for all }}x\in H.$ (3)
Immediately, we can see that the Hamiltonian $\mathcal{H}[u]$ can potentially be expressed as an inner product $\left\langle u,\kappa_{\theta}\right\rangle$ over the input function $u$ and a learnable function $\kappa_{\theta}$. This changes the problem of approximating a functional to the problem of approximating a function, which neural networks can readily achieve through universal approximation theorems [Pinkus_1999, LESHNO1993861].
A common inner product for arbitrary functions $u,v$ in a Hilbert space $H$ involves an integral: $\left\langle u,v\right\rangle=\int_{\Omega}u(x)v(x)\text{d}x$. One can check that this satisfies the basic properties of inner products and is also valid for vector-valued functions $u,v$. Based on this formulation, we propose to parameterize Neural Functionals by the Integral Kernel Functional (IKF):
$\mathcal{H}_{\theta}[u]=\int_{\Omega}\kappa_{\theta}(x)u(x)\text{d}x$ (4)
In this form, we observe a connection between the integral kernel functional and the integral kernel operator that neural operators are built on [kovachki_neuraloperator]. Specifically, the integral kernel operator is given by:
$\mathcal{K}_{\theta}(u)(x)=\int_{\Omega}\kappa_{\theta}(x,y)u(y)\text{d}y$ (5)
where a learnable operator $\mathcal{K}_{\theta}$ acting on an input function $u(x)\in\mathcal{U}$ produces an output function $\mathcal{K}(v)(x)\in\mathcal{V}$ for Banach spaces $\mathcal{U},\mathcal{V}$; this expression is equivalent to the IKF when evaluated at a single point $x$. We can make a similar modification to neural operators where the kernel is allowed to be nonlinear by giving function values as input: $\kappa_{\theta}(x,y,u(x),u(y))$. In our setting, we can adopt a nonlinear kernel by allowing the kernel to change based on the input function: $\kappa_{\theta}(x,u(x))$.
[…]
Implementation
There are two main implementation choices: approximating the integral and parameterizing the kernel. The integral can be approximated by a Riemann sum:
[…]
with quadrature weights $\mu_{i}$. Interestingly, for neural functionals the full Riemann sum can be computed since the sum is only calculated once; in neural operator literature, approximating the integral with a full Riemann sum is usually too costly and the sum is truncated [li2020neuraloperatorgraphkernel, multipole, li2023geometryinformedneuraloperatorlargescale] or represented in Fourier space through the convolution theorem [li2021fourierneuraloperatorparametric]. In our implementation, we evaluate different quadratures but find that the trapezoidal rule is enough to accurately estimate the integral.
To parameterize the kernel, we take the perspective from the Riesz representation theorem. In the linear case, the kernel is simply a function $\kappa_{\theta}(x)\in H$ that maps input coordinates to values in the codomain. This is exactly what a neural field is; therefore, we can inherit the substantial prior work on neural fields [xie2022neuralfieldsvisualcomputing] to effectively parameterize $\kappa_{\theta}$.
One useful concept is the conditional neural field. In particular, the nonlinear kernel $\kappa_{\theta}(x,u(x))$ can be seen as a conditional neural field whereby the kernel output changes based on the input function $u(x)$. This concept can also be further extended to distinguish between local and global conditioning [peng2020convolutionaloccupancynetworks]; in neural field literature, the field can be conditioned on local information $u_{i}=u(x_{i})$ or global information $\mathbf{u}=[u(x_{0}),\ldots,u(x_{n})]$. Conditioning mechanisms can also be implemented in various ways. In the simplest case, the conditional information $u_{i}$ or $\mathbf{u}$ is concatenated to the input coordinate $x$ [mehta2021modulatedperiodicactivationsgeneralizable, park2019deepsdflearningcontinuoussigned, mescheder2019occupancynetworkslearning3d], however, we adopt a approach based on Feature-wise Linear Modulation (FiLM) [perez2017filmvisualreasoninggeneral].
Beyond conditioning mechanisms, an important consideration is the architecture of the kernel itself. A powerful neural field architecture is the sinusoidal representation network or SIREN [sitzmann2020implicitneuralrepresentationsperiodic]. SIRENs have shown good performance not only in fitting a neural field $\kappa_{\theta}$, but also in accurately representing gradients $\nabla\kappa_{\theta}$, which can aid in fitting the functional derivative $\delta\mathcal{H}/\delta u$. With these architectural choices, a layer of the kernel $\kappa^{(l)}_{\theta}$ with local conditioning can be written as:
$\kappa^{(l)}_{\theta}(x_{i},u_{i})=\gamma^{(l)}_{\theta}(u_{i})\text{sin}(Wx_{i}+b)+\beta^{(l)}_{\theta}(u_{i})$ (7)
where $\gamma_{\theta}^{(l)}$ and $\beta_{\theta}^{(l)}$ are the FiLM scale and shift networks at layer $l$. Each SIREN layer has its own FiLM network and each FiLM network can also have multiple layers. Inputs $x_{i},u_{i}$ can optionally be lifted and the output $z_{i}$ can be projected back to the function dimension. Given the hierarchical nature of the proposed implementation, we also illustrate the architecture in Figure 1.
Figure 1: Left: The functional $\mathcal{H}[u]$ is approximated by an integral kernel functional (IKF). The integral is further approximated by a Riemann sum with a learnable kernel $\kappa_{\theta}$. Right: The kernel can be parameterized with various architectures and conditioning mechanisms. A local, FiLM-conditioned SIREN kernel is shown, where each SIREN layer has a dedicated FiLM network to produce scale and shift parameters. The output $z_{i}$ is for a single point $x_{i}$, which needs to be calculated $\forall i=\{1,\ldots,n\}$
A final consideration during implementation is batching. The layer update in Equation 7 is written in pointwise form, but we can observe that the computation does not depend on other points $\{(x_{j},u_{j}):j\neq i\}$, and the weights are shared between points. Therefore, the forward and backward pass for each point can be done in parallel; in practice, the spatial dimension is reshaped into the batch dimension for efficient training and inference.
Hamiltonian Neural Functional
So far, the developed architecture is largely agnostic to the downstream task, only requiring that the input function $u$ be discretized at a set of points $(x_{i},u_{i})$. To use neural functionals as a PDE surrogate model, we introduce the Hamiltonian Neural Functional (HNF). The time evolution of PDEs can be described by Hamilton’s equations, which depend on the operator $\mathcal{J}$ and the functional derivative $\delta\mathcal{H}/\delta u$. In Hamiltonian PDE systems, $\mathcal{J}$ is analytically derived or can be looked up, is guaranteed to be linear, and is usually simple. A common example in 1D is $\mathcal{J}=\partial_{x}$. Therefore, we choose to approximate $\mathcal{J}$ with a finite difference scheme, although it can also be learned with a neural operator. Lastly, we find that $\delta\mathcal{H}/\delta u$ can be computed directly with automatic differentiation. This is empirically shown in Section 4.1, but one way to see this is by considering the linear IKF. The gradient $\nabla_{\mathbf{u}}\sum_{i=1}^{n}\kappa_{\theta}(x_{i})u_{i}\>\mu_{i}\Delta x=[\kappa_{\theta}(x_{1}),\ldots,\kappa_{\theta}(x_{n})]$, up to a constant factor, which is the discretization of the learned function $\kappa_{\theta}(x)$. IN this case, the functional derivative is approximated by an arbitrary function, which is the desired behavior.
Algorithm 1 Training a HNF
1:repeat
2:  $\mathcal{H}_{\theta}\leftarrow\sum_{i=1}^{n}\kappa_{\theta}(x_{i},u_{i})u_{i}\>\mu_{i}\Delta x$
3:  $\frac{\delta\mathcal{H}_{\theta}}{\delta u}\leftarrow\text{autograd}(\mathcal{H}_{\theta},\mathbf{u})$
4:  $\mathcal{L}=||\frac{\delta\mathcal{H}_{\theta}}{\delta u}-\frac{\delta H}{\delta u}||^{2}\>\text{or}\>||\mathcal{J}(\frac{\delta\mathcal{H}_{\theta}}{\delta u})-\frac{\text{d}\mathbf{u}}{\text{d}t}||^{2}$
5:  $\theta\leftarrow\text{Update}(\theta,\nabla_{\theta}\mathcal{L})$
6:until converged
Within this framework, the training is described in Algorithm 1. For Hamiltonian systems, we assume knowledge of the analytical form of $\mathcal{H}$ and $\delta\mathcal{H}/\delta u$, and labels for $\delta\mathcal{H}/\delta u$ or $\text{d}\mathbf{u}/\text{d}t$ can be calculated from training data using finite difference schemes. The training loss can either be evaluated in the functional form $\delta\mathcal{H}_{\theta}/\delta u$ or the temporal form $\mathcal{J}(\delta\mathcal{H}_{\theta}/\delta u)$; we implement the former since it more directly optimizes $\kappa_{\theta}$.
During inference, the current state $\mathbf{u}^{t}$ and coordinates $\mathbf{x}$ are used in a forward pass to compute $\mathcal{H}_{\theta}$. A backward pass is then used to calculate $\delta\mathcal{H}_{\theta}/\delta u$; using the operator $\mathcal{J}$, a prediction for $\frac{\text{d}\mathbf{u}}{\text{d}t}|_{t=t}=\mathcal{J}(\frac{\delta\mathcal{H}_{\theta}}{\delta u})$ is calculated. The estimated derivative $\frac{\text{d}\mathbf{u}}{\text{d}t}$ is used to update the state $\mathbf{u}^{t+1}=\text{ODEint}(\mathbf{u}^{t},\frac{\text{d}\mathbf{u}}{\text{d}t})$ using a numerical integrator. In our implementation, we use a 2nd-order Adams-Bashforth method.
4 Experiments
4.1 Toy Examples
To better understand and situate neural functionals, we examine its ability to model simple, analytically constructed functionals and compare its performance to other architectures. The experiment setup is as follows: given a functional $\mathcal{F}[u]$, random polynomials $u(x)=c_{0}x^{p}+c_{1}x^{p-1}+\ldots+c_{p-1}x+c_{p}$ are generated by uniformly sampling $\{c_{i}\in[a,b]:i=0,\ldots,p\}$. Therefore, the dataset consists of $N$ pairs of polynomials and evaluated functionals $(u^{n}(x),\mathcal{F}[u^{n}(x)])$ for $n=1,\ldots,N$. In practice, each function $u^{n}(x)$ is discretized at a set of points $\{x_{i}:i=1,\ldots,M\}$, such that it is represented as $\mathbf{u}^{n}=[u^{n}(x_{1}),\ldots,u^{n}(x_{L})]$, and $\mathcal{F}[\mathbf{u}^{n}]$ remains a scalar.
We consider two cases: a linear functional $\mathcal{F}_{l}[u]=\int_{x_{1}}^{x_{M}}u(x)*x^{2}\text{d}x$ and a nonlinear functional $\mathcal{F}_{nl}[u]=\int_{x_{1}}^{x_{M}}(u(x))^{3}\text{d}x$. In both cases, constructed polynomials $u^{n}(x)$ can be substituted and the definite integral can be analytically solved to use as training labels. Additionally, the functional derivatives are $\frac{\delta\mathcal{F}_{l}}{\delta u}=x^{2}$ and $\frac{\delta\mathcal{F}_{nl}}{\delta u}=3u^{2}$, which are used to evaluate the gradient performance of models. We restrict the degree of $u^{n}(x)$ to $p=2$ and sample $c_{i}\in[-1,1]$; 100 training and 10 validation samples are generated for each case. Coordinates $x_{i}$ are uniformly spaced on the domain $[-1,1]$ with $M=100$ points.
Two baselines are considered: (1) a MLP that takes concatenated inputs $[\mathbf{x},\mathbf{u}]\in\mathbb{R}^{2M}$ and projects to an output $\mathcal{F}_{\theta}\in\mathbb{R}$ through hidden layers, and (2) a FNO with mean-pooling at the final layer to evaluate an operator baseline. The FNO baseline takes inputs with spatial and channel dimensions, or stacked inputs $[\mathbf{x},\mathbf{u}]\in\mathbb{R}^{M\times 2}$ and projects to an output field $\mathbf{z}\in\mathbb{R}^{M}$, which is then pooled to $\mathcal{F}_{\theta}\in\mathbb{R}$. For simplicity, the proposed neural functional (NF) is used with a local MLP kernel and is linear for $\mathcal{F}_{l}$ and nonlinear for $\mathcal{F}_{nl}$. In the nonlinear case, FiLM is not used and $[x_{i},u_{i}]\in\mathbb{R}^{2}$ is concatenated as input to $\kappa_{\theta}$. Lastly, models are trained with the loss $\mathcal{L}=\sum_{i=1}^{n}||\mathcal{F}[u^{n}(x)]-\mathcal{F}_{\theta}(\mathbf{u}^{n},\mathbf{x})||_{2}^{2}$; training is supervised with the functional to evaluate if models can implicitly learn accurate derivatives.
Base OOD ($c_{i}\in[1,3]$) Disc. ($x_{i}\in[-2,2]$)
Metric MLP FNO NF MLP FNO NF MLP FNO NF
$\mathcal{F}_{l}[u]$ 1.0e-5 5.4e-4 9.8e-16 0.043 0.097 2.7e-14 31.21 31.53 0.21
$\delta\mathcal{F}_{l}/\delta u$ 0.081 0.24 7.0e-4 0.14 0.29 7.0e-4 2.64 2.68 0.033
$\mathcal{F}_{nl}[u]$ 0.12 0.026 0.0023 6131 5864 2126 315.9 281.0 62.3
$\delta\mathcal{F}_{nl}/\delta u$ 1.99 1.84 0.089 1684 1659 998 74.6 69.9 29.5
Table 1: Validation MSE is reported across different experiments with linear and nonlinear functionals. Errors are calculated in the scalar $\mathcal{F}[u]$ and gradient domains $\delta\mathcal{F}/\delta u$ to evaluate performance in fitting functionals and implicitly learning functional derivatives. Additionally, generalization to OOD inputs and unseen discretizations are evaluated. Parameter counts are: MLP (7.5K), FNO (10.9K) and NF (4.3K); each model takes roughly 30s to train until convergence.
To further evaluate generalization, after training we consider two cases: (1) OOD, where validation inputs $u(x)$ are sampled with polynomial coefficients $c_{i}\in[1,3]$ to model out-of-distribution functions, (2) Disc., where validation inputs $u(x)$ are uniformly discretized on a larger domain $x_{i}\in[-2,2]$ with a larger $\Delta x$, while keeping $M=100$. In each experiment, validation errors are reported with the metric $\frac{1}{n}\sum_{i=1}^{n}||\mathcal{F}[u^{n}]-\mathcal{F}_{\theta}(\mathbf{u}^{n},\mathbf{x})||_{2}^{2}$ or $\frac{1}{n}\sum_{i=1}^{n}||\frac{\delta\mathcal{F}}{\delta u^{n}}-\text{autograd}(F_{\theta},\mathbf{u}^{n})||_{2}^{2}$ to measure performance in fitting functionals and learning accurate functional derivatives.
Results are given in Table 1. We observe that neural functionals can accurately represent functionals and their derivatives, outperforming other architectures. Despite only training on scalars $\mathcal{F}[u]$, neural functionals can implicitly learn the correct functional derivatives, as plotted in Figure 2. This is new; even on toy problems, traditional architectures do not have well-behaved functional derivatives, as such, learning in the gradient domain for Hamiltonian systems would not be possible. Although the gradient performance is poor, conventional architectures can still fit functionals and the MSE for validation samples within the training distribution (Base) is low. When considering generalization performance, neural functionals can readily generalize to unseen input functions as well as unseen discretizations in the linear case. This empirically observed generalization is theoretically motivated; the Riesz representation theorem guarantees that a single kernel $\kappa_{\theta}$ can represent a linear functional $\mathcal{F}_{l}$ for all possible input functions $u(x)$ (e.g., polynomial, exponential, etc.). Generalization performance with nonlinear functionals is more challenging, as larger polynomial coefficients $c_{i}$ and larger coordinates $x_{i}$ can quickly increase $F_{nl}[u(x)]$ due to large exponents. Despite this, neural functionals still outperform other architectures in these domains, and learn accurate nonlinear functionals and their derivatives for samples within distribution.
Figure 2: After training to regress $\mathcal{F}[u]$, the gradients of the network with respect to an input function $u$ are plotted for different models. Left, A: Neural functionals are able to accurately model derivatives, even on unseen discretizations, whereas other architectures fail even on their training distribution. Right, B: Nonlinear functionals have more complex derivatives that change with input functions $u$. In this instance, a quartic function is learned and the model extrapolates to unseen discretizations.
4.2 1D Advection and KdV Equations
Formulation
To evaluate the Hamiltonian Neural Functional (HNF) on PDE problems, we consider the 1D Advection and the 1D Korteweg–De Vries (KdV) equation. The Hamiltonians for each PDE are given as [miura_kdv, kdv_1]:
$\mathcal{H}_{adv}[u]=\int_{\Omega}-\frac{1}{2}(u(x))^{2}\text{d}x,\qquad\mathcal{H}_{kdv}[u]=\int_{\Omega}-\frac{1}{6}(u(x))^{3}-u(x)\frac{\partial^{2}u}{\partial x^{2}}(x)\text{d}x,$ (8)
Given an initial state $u^{0}$, the Hamiltonian for each PDE is conserved over time (i.e., $\mathcal{H}[u^{t}]=\mathcal{H}[u^{0}]$). For both PDEs, the operator $\mathcal{J}$ is defined as $\partial_{x}$. We can check that Hamilton’s equations hold for the 1D Advection equation:
$\frac{\delta\mathcal{H}_{adv}}{\delta u}=-u(x),\qquad\mathcal{J}(\frac{\delta\mathcal{H}_{adv}}{\delta u})=-\frac{\partial u}{\partial x}=\frac{\text{d}u}{\text{d}t}$ (9)
which recovers the correct PDE for 1D Advection. A similar calculation can be done to recover the PDE for the 1D KdV equation ($u_{t}+uu_{x}+u_{xxx}=0$) from Hamilton’s equations by using $\frac{\delta\mathcal{H}_{kdv}}{\delta u}=-\frac{1}{2}u^{2}-u_{xx}$. Therefore, given a training dataset of spatiotemporal PDE data, the labels $\frac{\delta\mathcal{H}}{\delta u}$ can be computed using finite differences at every timestep $t$. The gradients of the HNF network $\nabla_{\mathbf{u}}(\mathcal{H}_{\theta}(\mathbf{x},\mathbf{u}^{t}))$ are then fitted to these labels.
Evaluation
Data for the 1D Advection and KdV equations are generated from a numerical solver using random initial conditions [zhou2024maskedautoencoderspdelearners, brandstetter2022liepointsymmetrydata]:
$u^{0}(x)=\sum_{i=1}^{J}A_{i}*\text{sin}(\frac{2\pi l_{i}*x}{L}+\phi_{i})$ (10)
where $L$ is the length of the domain, $A_{i}\in[-0.5,0.5]$, $l_{i}\in\{1,2,3\}$, and $\phi_{i}\in[0,2\pi]$; $J=5$ for Advection and $J=10$ for KdV. The Advection equation is solved on a uniform domain $x_{i}\in[0,16]$ with $n_{x}=128$ and the KdV equation is solved on a domain $x_{i}\in[0,100]$ with $n_{x}=256$. Both equations have periodic boundary conditions. For the Advection equation, 1024/256 samples are generated for train/validation, and for the KdV equation, 2048/256 samples are generated. To evaluate temporal extrapolation, the training dataset is truncated to a shorter time horizon. For the Advection equation, it is solved from $t=0s$ to $t=4s$ during training ($n_{t}=200$) and evaluated from $t=0s$ to $t=20s$ during validation ($n_{t}=1000$); a similar scenario is constructed for the KdV equation with $t=0s$ to $t=25s$ for training ($n_{t}=50$) and $t=0s$ to $t=100s$ for validation ($n_{t}=200$).
Adv KdV
Metric: Roll. Err. $\downarrow$ Corr. Time $\uparrow$
FNO 0.81$\pm$0.17 69.5$\pm$8.2
Unet 0.52$\pm$0.35 125.7$\pm$8.5
FNO($\frac{d\mathbf{u}}{dt}$) 0.044$\pm$0.002 75.5$\pm$1.9
Unet($\frac{d\mathbf{u}}{dt}$) 0.068$\pm$0.029 127.4$\pm$4.9
HNF 0.0039$\pm$0.0002 150.9$\pm$3.3
Table 2: Results for 1D PDEs. Parameter counts are: FNO (65K), Unet (65K), HNF (32K) for Adv, and FNO (135K), Unet (146K), HNF (87K) for KdV.
Figure 3: Hamiltonians for Adv and KdV for predicted trajectories, plotted over time. FNO and Unet models are shown with their $d\mathbf{u}/dt$ variants. HNFs can better conserve Hamiltonians and remain stable over time.
Two baselines are considered, a FNO [li2021fourierneuraloperatorparametric] and a Unet [gupta2022multispatiotemporalscalegeneralizedpdemodeling] model. Additionally, the HNF model is used with a nonlinear, FiLM-conditioned SIREN kernel. Global conditioning is used in the KdV experiments by calculating the scale/shift parameters as $\gamma_{i}^{(l)}=[\gamma_{\theta}^{(l)}(\mathbf{u}^{t})]_{i}$ and $\beta_{i}^{(l)}=[\beta_{\theta}^{(l)}(\mathbf{u}^{t})]_{i}$; the global FiLM networks are each parameterized by a shallow, 1D CNN. We find that global conditioning is needed in more complex PDEs since the kernel output $z_{i}=k_{\theta}(\cdot)$ can depend on more than just $u_{i}$ at a given point $x_{i}$. In particular, the Hamiltonian for the KdV equation has a $u_{xx}$ term that requires non-local knowledge of $\mathbf{u}^{t}$. Lastly, to ensure a fair comparison, we also train variants of the FNO and Unet baselines that predict temporal derivatives $\text{d}\mathbf{u}/\text{d}t$ [Zhou_2025], as opposed to the usual variants that predict $\mathbf{u}^{t+1}$; these use the same numerical integrator during inference as the HNF model.
The metrics that are reported are the rollout error = $\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}(\mathbf{u}^{t},\mathbf{u}_{\theta}^{t})$, and the correlation time = $\max(t)\>s.t.\>\mathcal{C}(\mathbf{u}^{t},\mathbf{u}_{\theta}^{t})>0.8$. The loss used is a relative L2 loss, or $\mathcal{L}(\mathbf{u}^{t},\mathbf{u}_{\theta}^{t})=\frac{||\mathbf{u}^{t}-\mathbf{u}^{t}_{\theta}||_{2}^{2}}{||\mathbf{u}^{t}||_{2}^{2}}$ and the correlation criterion $\mathcal{C}$ is the Pearson correlation. Rollout error is used to measure the accuracy of the predicted trajectory; for chaotic dynamics rollout error is often skewed after autoregressive drift, therefore correlation time measures the portion of the trajectory that is accurate. Each metric is averaged over the validation set and the mean and standard deviation over three seeds is reported in Table 2. Furthermore, given a predicted trajectory $[\mathbf{u}_{\theta}^{1},\ldots,\mathbf{u}_{\theta}^{T}]$ from a model, we can examine its Hamiltonian at each timestep and plot the scalar value $\mathcal{H}[\mathbf{u}_{\theta}^{t}]$ over time, shown in Figure 3.
We observe that the proposed HNF model is able to outperform current baselines, both on quantitative metrics and when conserving Hamiltonians over time. HNFs are also efficient, capable of achieving better performance with nearly half the parameters of baseline models due to sharing weights between input points $x_{i},u_{i}$. Furthermore, despite training on a shorter time horizon, HNFs can extrapolate and remain stable in longer prediction horizons. This is more pronounced in the KdV equation, where initial states are smoother than later, more chaotic states. An example trajectory is shown in Figure 5, where the baselines struggle to predict later states since these are not seen during training.
4.3 2D Shallow Water Equations
Formulation
To evaluate HNFs on a more complex system, we consider the 2D Shallow Water Equations (SWE). These equations are an approximation of the Navier-Stokes equations when the horizontal length scale is much larger than the vertical length scale. We consider its conservative form, which is often used as a simple model for atmosphere or ocean dynamics.
For vector-valued functions $\mathbf{u}(x,y)\in\mathbb{R}^{d}$, the operator $\mathcal{J}$ becomes a $d\times d$ matrix of operators. For the 2D Shallow Water Equations, the operator matrix $\mathcal{J}$ is given by [ALLEN1996229, Ripa1993, Dellar2003]:
$\mathcal{J}=\begin{bmatrix}0&-q&\partial_{x}\\ q&0&\partial_{y}\\ \partial_{x}&\partial_{y}&0\end{bmatrix},\qquad q=\frac{\partial v_{x}}{\partial y}-\frac{\partial v_{y}}{\partial x}$ (11)
where $q$ is the vorticity. The Hamiltonian for this system is:
$\mathcal{H}[\mathbf{u}]=\mathcal{H}[v_{x},v_{y},h]=\int_{\Omega}\frac{1}{2}h(v_{x}^{2}+v_{y}^{2})+\frac{1}{2}gh^{2}\text{d}A$ (12)
where $g$ is the gravitational constant and the velocities and height $v_{x},v_{y},h$ are usually defined on a grid (i.e., $v_{x}=v_{x}(x_{i},y_{j})$). One can verify Hamilton’s equations and that $\mathcal{J}\frac{\delta\mathcal{H}}{\delta\mathbf{u}}$ recovers the 2D Shallow Water Equations:
$\partial_{t}h+\nabla\cdot(\mathbf{v}h)=0,\qquad\partial_{t}\mathbf{v}+\mathbf{v}\cdot\nabla\mathbf{v}=-g\nabla h$ (13)
Evaluation
Data for the 2D SWE system is generated from PyClaw [clawpack, pyclaw-sisc, mandli2016clawpack] using random sinusoidal initial conditions (Sines) for the height:
$h^{0}(x,y)=H+\sum_{i=1}^{J}A_{i}*\text{sin}(\frac{2\pi m_{i}*x}{L_{x}}+\frac{2\pi n_{i}*y}{L_{y}}+\phi_{i})$ (14)
with initial velocities set to zero. Constants are randomly sampled from $A_{i}\in[-0.1,0.1]$, $m_{i},n_{i}\in\{1,2,3\}$, and $\phi_{i}\in[0,2\pi]$ with $J=5$, $H=1$, and $L_{x}=L_{y}=2$. The domain is $(x_{i},y_{i})\in[-1,1]^{2}$ with $n_{x}=n_{y}=128$; each sample has $n_{t}=100$ timesteps from $t=0s$ to $t=2s$. The gravitational constant is set to $g=1$; 256 samples are generated for training and 64 samples are generated for validation. Periodic boundary conditions are used.
Model Sines Pulse
FNO 0.057$\pm$0.002 0.117$\pm$0.0009
Unet 0.010$\pm$0.0014 0.042$\pm$0.0006
HNF 0.026$\pm$0.0003 0.021$\pm$0.0015
Table 3: Rollout errors for 2D SWE. Parameter counts are: FNO (7M), Unet (3M), HNF (3M). After training, models are evaluated on in-distribution ICs (Sines) or out-of-distribution ICs (Pulse).
Figure 4: Hamiltonians for Sines and Pulse ICs, plotted over time. Despite potentially larger rollout errors, HNFs can better conserve the Hamiltonian of predicted solutions.
To evaluate model generalization, we generate an additional test set of 64 samples with a random Gaussian pulse as initial conditions (Pulse). The initial height is set to an isotropic Gaussian centered at the origin with random covariance $\sigma\in[0.1,0.5]$ and the height is scaled to be between $[1,1.5]$; initial velocities are zero. The domain, BCs, and discretizations remain the same.
We again consider benchmarking against Unet and FNO models for 2D systems, and use a global, FiLM-conditioned SIREN kernel for the HNF model. Validation rollout errors for in-distribution ICs (Sines) and out-of-distribution ICs (Pulse) are given in Table 3, where the mean and standard deviation is calculated over three seeds. Additionally, the Hamiltonian over time for model rollouts on these test sets are plotted in Figure 4. Interestingly, the derivative variants of the baseline models could not train stably, possibly due to sharp gradients that develop in the absence of viscosity. Additionally, the baselines are trained with the benefit of normalization, as the height channel has a different scale than the velocity channels; however, since the Hamiltonian is not invariant under scaling and shifting, normalization was not performed for the HNF model.
For test samples within the training distribution (Sines), conventional neural surrogates are able to perform well and achieve low loss, and HNFs are on-par with these models. However, when examining the Hamiltonian, HNFs can predict solutions that better conserve physical invariances. We validate this in Appendix B.1 by calculating the error of the Hamiltonian of predicted trajectories across different experiments. This ability to conserve the Hamiltonian helps HNFs to generalize to unseen initial conditions, and it achieves better performance on the Gaussian Pulse test set. Interestingly, it also conserves the Hamiltonian on out-of-distribution inputs, suggesting a strong inductive bias in the model. This is an empirically observed bias, to understand what mechanisms enable energy conservation, we ablate different methods within HNFs to understand their contributions, shown in Appendix B.1. Lastly, to visualize model performance, predicted trajectories of Sines and Pulse ICs are shown in Figures 6 and 7.
5 Discussion
Limitations
Despite an interesting theoretical basis and its novelty, HNFs have some limitations that we hope future work can address. Firstly, the theory for nonlinear functionals is underdeveloped. Another limitation is the additional overhead during inference to backpropagate through the network and evaluate the linear operator $\mathcal{J}$. We quantify this in section B.2 in the Appendix. Additionally, implementing $\mathcal{J}$ without a neural operator may require care to avoid numerical errors; this is further discussed in Appendix D.1. Furthermore, HNFs are only applicable to Hamiltonian systems; this would exclude systems with viscosity or friction, although modifications exist to extend Hamiltonian mechanics to non-conservative systems [sosanya2022dissipativehamiltonianneuralnetworks, hh_decomp]. Even for conservative PDE systems, implementing a Hamiltonian structure is not as straightforward as training a neural surrogate with next-step prediction. Lastly, while not a limitation, an interesting observation is that introducing a loss on $\mathcal{H}[\mathbf{u}^{t}]-\mathcal{H}[\mathbf{u}_{\theta}^{t}]$ does not improve HNF performance; we hypothesize that since the Hamiltonian is conserved over time, learning the identity mapping is the easiest way to minimize this error and thus degrades model training.
Outlook
Despite these limitations, neural functionals represent a new class of models for approximating functionals and their derivatives. Within a Hamiltonian framework, HNFs show good performance in predicting PDE solutions and have excellent capabilities in conserving energy-like quantities. Interestingly, molecular dynamics (MD) is a highly promising use case; on the atomistic scale, energies are always conserved and its derivative (forces) are used to evolve simulations. For ML-based MD simulations, conserving energy is highly important due to extremely long rollouts, as timesteps $\Delta t$ are on the picosecond scale [fu2023forcesenoughbenchmarkcritical]. Apart from MD, functionals arise naturally in many physical systems. While the applications are still speculative, functionals are a concise and meaningful description of a physical state, such as describing the drag or weight of a continuum system, and having meaningful functional derivatives may have interesting uses for optimization.
6 Related Works
To the best of the authors’ knowledge, this is the first time an architecture has been developed to learn function to scalar mappings. Therefore, we focus on discussing related works in PDE surrogate modeling and neural networks with Hamiltonian structure. Original work on Hamiltonian Neural Networks (HNNs) [greydanus2019hamiltonianneuralnetworks, onlearninghamiltonian] established them as a promising architecture to respect physical laws, while later work improved their performance [DAVID2023112495, HORN2025113536, Zhong2020Symplectic, Chen2020Symplectic, sanchezgonzalez2019hamiltoniangraphnetworksode], investigated the mechanisms behind HNNs [gruver2022deconstructinginductivebiaseshamiltonian], or used HNNs in different applications [Mattheakis_2022, Toth2020Hamiltonian, zhong2021extendinglagrangianhamiltonianneural]. A closely related set of works develop neural networks in a Lagrangian framework [cranmer2019lagrangian, lutter2019deeplagrangiannetworksusing], which also preserve learned energies. More broadly, learning an energy and optimizing its gradient is an established framework in ML-based molecular dynamics simulation [GDML, gasteiger2022directionalmessagepassingmolecular, schütt2017schnetcontinuousfilterconvolutionalneural, zhang2018endtoendsymmetrypreservinginteratomic, Batzner_2022]. Despite research in this area, no prior works have considered learning with infinite-dimensional Hamiltonian systems and most HNNs have only been demonstrated in simple, particle-based systems.
Neural PDE surrogates are a growing field, with many works proposing architectures to improve model accuracy [lippe2023pderefinerachievingaccuratelong, wu2024transolverfasttransformersolver, alkin2025universalphysicstransformersframework], generalizability [zhou2024strategiespretrainingneuraloperators, hao2024dpotautoregressivedenoisingoperator, herde2024poseidonefficientfoundationmodels], or stability [zhou2025text2pdelatentdiffusionmodels, li2022learningdissipativedynamicschaotic, LIST2025117441, li2025generativelatentneuralpde]. Specific to our work, many PDE surrogates have also employed physics-based prior knowledge to improve model performance and adherence to governing laws. The most straightforward approach is to use physics-informed losses [RAISSI2019686, li2023physicsinformedneuraloperatorlearning, wang2021learningsolutionoperatorparametric], but another, more fundamental approach is to enforce physical symmetries and equivariances [shumaylov2025liealgebracanonicalizationequivariant, wang2021incorporating, brandstetter2023cliffordneurallayerspde, akhoundsadegh2023liepointsymmetryphysics, horie2023physicsembeddedneuralnetworksgraph]. Interestingly, by Noether’s theorem, every symmetry has an associated conservation law, which HNFs rely on as an inductive bias; through this lens the current work is another way of softly enforcing symmetries in predicted solutions, specifically time-translation symmetry. Despite previous works that incorporate physical priors into models, this work is likely the first to extend Hamiltonian mechanics to PDE surrogates, enabled by introducing neural functionals.
7 Conclusion
In this work, we have proposed a novel method for learning function to scalar mappings, enabling the extension of Hamiltonian mechanics to neural PDE surrogates. We demonstrate that neural functionals are able to implicitly learn functional derivatives, as well as achieve good performance on predicting 1D and 2D PDE solutions. In addition, Hamiltonian Neural Functionals have good capabilities in conserving energy-like quantities, enabling it to generalize to certain out-of-distribution inputs, such as a longer time horizon or unseen initial conditions. We hope that future work can continue to improve neural functionals as well as take advantage of functional approximation in new and insightful applications.
Acknowledgments and Disclosure of Funding
[…]
Appendix A Additional Visualizations
Figure 5: Predicted trajectories of the KdV equation. The white dashed line denotes the training horizon, where only states from $t=0$ to $t=50$ are seen during training. Despite not seeing later, more chaotic states, HNFs still extrapolate beyond the training horizon by conserving the Hamiltonian, while other models struggle. FNO and Unet models are shown with their $d\mathbf{u}/dt$ variants.
Figure 6: Predicted trajectories of the Shallow Water Equations (SWE) with sinusoidal ICs, displayed at $t=10$ and $t=50$. All models are able to achieve good performance on in-distribution samples.
Figure 7: Predicted SWE trajectories with Gaussian pulse ICs, displayed at $t=10$ and $t=50$. Despite showing different behavior, HNFs can better generalize to OOD samples.
Appendix B Additional Results
B.1 Hamiltonian Errors and Inductive Bias
Model Adv KdV SWE-Sines SWE-Pulse
FNO 2.18 NaN 0.0091 0.1326
Unet 0.22 2.37 0.0053 0.0158
FNO($\frac{d\mathbf{u}}{dt}$) 0.033 6.75e8 - -
Unet($\frac{d\mathbf{u}}{dt}$) 0.043 5.76e6 - -
HNF 0.0002 1.32 0.0015 0.0003
Table 4: Relative L2 Error for each experiment, evaluated on the Hamiltonian of predicted trajectories. Despite not including the Hamiltonian in the training loss and testing on OOD samples, HNFs are exceptional at predicting solutions that conserve the Hamiltonian.
In the main body, we presented qualitative evidence that HNFs are able to better conserve the Hamiltonian of predicted trajectories. To quantify this claim, we calculate the Relative L2 Error of the Hamiltonian of predicted trajectories across all PDE experiments. Specifically, the error of a sample is given by: $\frac{1}{T}\sum_{t=1}^{T}\frac{||\mathcal{H}[\mathbf{u}^{t}]-\mathcal{H}[\mathbf{u}_{\theta}^{t}]||_{2}^{2}}{||\mathcal{H}[\mathbf{u}^{t}]||_{2}^{2}}$, where $\mathbf{u}^{t}$ is the true solution at time $t$, $\mathbf{u}^{t}_{\theta}$ is the predicted solution at time $t$, and $\mathcal{H}[\cdot]$ is the Hamiltonian given by the integrals in Equations 8 and 12. The Hamiltonian errors are averaged across the validation sets and reported in Table 4. We note that in the KdV experiments, the dynamics are chaotic and the Hamiltonian is highly nonlinear, both of which contribute to large errors once predicted trajectories diverge. Overall, we find that HNFs have exceptional capabilities in implicitly conserving energy-like quantities in their prediction.
[…]
  1. 1.
    ODE Bias: HNFs predict $\frac{d\mathbf{u}}{dt}$ and use an ODE integrator to evolve PDE dynamics.
  2. 2.
    Hamiltonian Bias: HNFs rely on $\mathcal{J}\frac{\delta\mathcal{H}}{\delta\mathbf{u}}$ to calculate $\frac{d\mathbf{u}}{dt}$.
  3. 3.
    Gradient Learning Bias: HNFs rely on $\text{autograd}(\mathcal{H}_{\theta}[\mathbf{u}],\mathbf{u})$ to calculate $\frac{\delta\mathcal{H}}{\delta\mathbf{u}}$.
  4. 4.
    Neural Functional Bias: HNFs rely on neural functionals to calculate $\text{autograd}(\mathcal{H}_{\theta}[\mathbf{u}],\mathbf{u})$.
We note that these biases are listed from simplest to most complex and are cumulative; for example, it is not possible to implement Hamiltonian structure without also implementing an ODE integrator. We modify a Unet to incorporate each inductive bias and evaluate its performance on the KdV equation, which is the most chaotic setting. The ODE bias was considered in the main work by training $d\mathbf{u}/dt$ model variants, while incorporating Hamiltonian bias is done by changing the training label to $\delta\mathcal{H}/\delta\mathbf{u}$ and using $\mathcal{J}$ during inference. To extend Unets to gradient learning, we add a linear head to project the output field to a scalar and use automatic differentiation to obtain $\delta\mathcal{H}/\delta\mathbf{u}$. After training, we evaluate model correlation time and Hamiltonian error on the validation set and the results are reported in Table 5.
We find that the main contributors to HNF performance and its ability to conserve energies over time is using Hamiltonian structure as well as the neural functional architecture. Using an ODE integrator alone degrades model performance and implementing gradient learning with Unets also harms performance. Interestingly, using gradient learning with conventional architectures disproportionately harms energy conservation; this implies that in PDEs, gradient domain learning improves energy conservation only when architectures have meaningful functional derivatives. This is not readily apparent, as observations from molecular dynamics or HNNs suggest that gradient domain learning generally improves energy conservation without considering the underlying architectures [greydanus2019hamiltonianneuralnetworks, cranmer2019lagrangian, GDML]. Based on these findings, we hypothesize the following mechanism underlying HNFs: Neural functionals enable the learning of functional derivatives, combining this with infinite-dimensional Hamiltonian mechanics leads to its performance and conservation properties. Conveniently, these two contributions are the main novelties of this work.
Metric: Correlation Time ($\uparrow$) Hamiltonian Error ($\downarrow$)
Unet (Base) 125.25 2.37
Unet (ODE) 120.5 5.76e6
Unet (Ham.) 143.75 2.06
Unet (Grad.) 141 4.71
HNF 151.75 1.32
Table 5: Correlation time and Hamiltonian errors for Unet models with increasingly more inductive biases, compared to HNFs on the KdV equation. Using ODE integrators or gradient domain learning both degrade performance, while using Hamiltonian structure or neural functionals both increase performance and energy conservation.
B.2 Timing Experiments
Model (#Params) Adv KdV SWE-Sines
FNO (65K/135K/7M) 0.0829 0.091 0.967
Unet (65K/146K/3M) 0.138 0.146 1.345
HNF (32K/87K/3M) 0.126 0.228 4.547
Table 6: Computational cost per inference step for each model across different PDEs, given in milliseconds (ms). Model sizes are also reported for the (Adv/Kdv/SWE) experiments. Each run uses a single NVIDIA RTX 6000 Ada GPU.
The additional overhead of using automatic differentiation, evaluating the operator $\mathcal{J}$, and applying an ODE integrator during inference increases the computational cost of HNFs. We examine this in Table 6 by performing a model rollout and averaging the time per inference step across a batch. Consistent with prior works, FNO models remain a fast neural PDE surrogate. While HNFs may have lower parameter counts, this may be a misleading measure of inference speed. Automatic differentiation calculates the gradient with respect to each input point, therefore the computational cost of this step can scale with the size of the input grid. This can be seen when comparing the Adv and KdV experiments, where the grid size doubles from $n_{x}=128$ to $n_{x}=256$, and the computational cost roughly doubles as well. In 2D experiments, the cost of HNFs is appreciably more than other baselines. We do not anticipate that the finite-difference (FD) schemes used in $\mathcal{J}$ or ODE integrators contribute meaningfully to computational cost, since FD schemes use $\mathcal{O}(M)$ operations, where $M$ is the number of grid points, and numerical integration occurs in constant time.
Appendix C Proof of Linear Functional Approximation
[…]
Let $D\subset\mathbb{R}^{d}$ be a domain. Let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ denote the natural numbers that include zero. Let $\mathcal{X}$ be a Banach space, and $\mathcal{X}^{*}$ be its continuous dual space. Specifically, the dual space contains all continuous, linear functionals $f:\mathcal{X}\rightarrow\mathbb{R}$ with the norm: $||f||_{\mathcal{X}^{*}}=\sup_{x\in\mathcal{X},||x||_{\mathcal{X}}=1}|f(x)|<\infty$.
For any multi-index $\alpha\in\mathbb{N}_{0}^{d}$, $\partial^{\alpha}$ is the $\alpha$-th partial derivative of $f$. The following spaces are defined for $m\in\mathbb{N}_{0}$:
  • $C(D)=\{f:D\rightarrow\mathbb{R}:f\text{ is continuous}\}$
[…]
  • $C_{c}^{\infty}(D)=\{f\in C^{\infty}(D):\text{supp}(f)\subset D\text{ is compact}\}$
  • $C^{m}_{b}(D)=\{f\in C^{m}(D):\max_{0\leq|\alpha|_{1}\leq m}\sup_{x\in D}|\partial^{\alpha}f(x)|<\infty\}$
  • $C^{m}(\bar{D})=\{f\in C^{m}_{b}(D):\partial^{\alpha}f\text{ is uniformly continuous }\forall 0\leq|\alpha|_{1}\leq m\}$
[…]
$N_{n}(\sigma;\mathbb{R}^{d},\mathbb{R}^{d^{\prime}}):=\{f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}:$ $f(x)=W_{n}\sigma(\ldots W_{1}\sigma(W_{0}x+b_{0})+b_{1}\ldots)+b_{n},$
$W_{0}\in\mathbb{R}^{d_{0}\times d},W_{1}\in\mathbb{R}^{d_{1}\times d_{0}},\ldots,W_{n}\in\mathbb{R}^{d^{\prime}\times d_{n-1}},$
$b_{0}\in\mathbb{R}^{d_{0}},b_{1}\in\mathbb{R}^{d_{1}},\ldots,b_{n}\in\mathbb{R}^{d^{\prime}},d_{0},d_{1},\ldots,d^{\prime}\in\mathbb{N}\}$
The activation functions $\sigma$ are restricted to the set:
$\sigma\in A_{m}:=\{\sigma\in C(\mathbb{R}):\exists n\in\mathbb{N}\>\text{s.t. }N_{n}\text{ is dense in }C^{m}(K)\forall K\subset\mathbb{R}^{d}\text{ compact}\}$
[…]
Let $D\subset\mathbb{R}^{d}$ be a domain and $L\in(C(\bar{D}))^{*}$ be a linear functional. For any compact set $K\subset C(\bar{D})$ and $\epsilon>0$, there exists a function $\kappa\in C^{\infty}_{c}(D)$ such that:
$\sup_{u\in K}|L(u)-\int_{D}\kappa u\text{d}x|<\epsilon$
This result establishes that a linear functional acting on functions in $C(\bar{D})$ can be approximated by a smooth integral kernel with arbitrary error $\epsilon$. To show functional approximation with a neural network, we slightly modify an existing lemma:
[…]
Let $D\subset\mathbb{R}^{d}$ be a domain and $\mathcal{A}=C(\bar{D})$. Let $L\in\mathcal{A}^{*}$ be a linear functional. For any compact set $K\subset\mathcal{A}$, $\sigma\in A_{0}$, and $\epsilon>0$, there exists a number $L\in\mathbb{N}$ and neural network $\kappa_{\theta}\in N_{L}(\sigma;\mathbb{R}^{d},\mathbb{R})$ such that:
$\sup_{u\in K}|L(u)-\int_{D}\kappa_{\theta}(x)u(x)\text{d}x|_{1}<\epsilon$
Proof.
Since $K$ is bounded, there exists a number $M>0$ such that:
$\sup_{u\in K}||u||_{\mathcal{A}}\leq M$
By Lemma 30, there exists a function $\kappa\in C_{c}^{\infty}(D)$ such that:
$\sup_{u\in K}|L(u)-\int_{D}\kappa u\text{d}x|<\frac{\epsilon}{2}$
Since $\sigma\in A_{0}$, there exists some $L\in\mathbb{N}$ and a neural network $\kappa_{\theta}\in N_{L}(\sigma;\mathbb{R}^{d},\mathbb{R})$ such that:
$||\kappa_{\theta}-\kappa||_{C}\leq\frac{\epsilon}{2M|D|}$
Then for any $u\in K$:
$|L(u)-\int_{D}\kappa_{\theta}(x)u(x)\text{d}x|_{1}$ $\leq|L(u)-\int_{D}\kappa u\text{d}x|_{1}+|\int_{D}(\kappa-\kappa_{\theta})u\text{d}x|_{1}$
$\leq\frac{\epsilon}{2}+M|D|*||\kappa-\kappa_{\theta}||_{C}$
$\leq\epsilon$
Therefore, linear functionals for continuously differentiable function classes can be approximated to arbitrary accuracy by a neural network using an integral kernel. Extensions to Sobolev and Hilbert spaces can also be proved in a similar manner.
[…]
Appendix D Implementation Details
D.1 Numerical Methods
Numerical Integration
ODE integration is performed using a 2nd-order Adams-Bashforth scheme. The solution at the next timestep $\mathbf{u}^{t+1}$ is calculated using the current timestep $\mathbf{u}^{t}$ and an estimate of the current derivative $\frac{d\mathbf{u}}{dt}|_{t=t}=f(\mathbf{\mathbf{u}}^{t})$:
$\mathbf{u}^{t+1}=\mathbf{u}^{t}+\Delta t\left(\frac{3}{2}f(\mathbf{\mathbf{u}}^{t})-\frac{1}{2}f(\mathbf{\mathbf{u}}^{t-1})\right)$ (15)
This can be implemented in constant time by caching prior derivative estimates $\{f(\mathbf{u}^{i}):i<t\}$, allowing the ODE integrator to have a 2nd-order truncation error ($\mathcal{O}(\Delta t^{2})$) while remaining fast. At the first timestep, there are no cached derivatives, therefore a first-order Euler integrator is used: $\mathbf{u}^{t+1}=\mathbf{u}^{t}+\Delta tf(\mathbf{\mathbf{u}}^{t})$. For the given experiments, these integrators are accurate since $\Delta t$ is small enough; to be used with larger timesteps $\Delta t$, higher-order integrators may be needed or smaller, intermediate steps will need to be taken [Zhou_2025]. This adds additional consideration to the training and deployment of derivative-based neural PDE surrogates.
Finite Difference Schemes
Implementing infinite-dimensional Hamiltonian structure involves approximating the operator $\mathcal{J}$, either through numerical methods or neural operators. We opt for numerical methods, as the operators are simple and in all our tests only involve first-order spatial derivatives $\partial_{x}$. For the 1D Advection equation, we implement a 2nd-order central difference scheme. For the 1D KdV equation, the spatial derivatives exhibit larger gradients, therefore we use an 8th-order central difference scheme with a smoothing stencil to approximate $\partial_{x}$ [snrd]. A similar scheme is implemented in 2D to approximate $\partial_{x}$ and $\partial_{y}$ for the 2D SWE experiments, and additionally, a Savitzky-Golay filter is used to further damp numerical oscillations. These arise since the lack of viscosity creates discontinuities in the shallow-water equations, alternative methods may include flux-limiting or non-oscillatory schemes [HARTEN19973, LIU1994200]. While these implementations work, they still required tuning, as such, the use of a neural operator to approximate $\mathcal{J}$ to make HNFs fully learnable would be an interesting future direction.
Additionally, to calculate temporal derivatives $d\mathbf{u}/dt$ during training, a 4-th order Richardson extrapolation is used to provide accurate derivatives. At the boundaries $t=0,\>t=T$ a one-sided Richardson extrapolation is used [RAHUL200613].
Numerical Solvers
[…]
D.2 Experimental Details
Hyperparameters
Each model has different hyperparameters for different PDEs (Adv/KdV/SWE). We provide key hyperparameters for these experiments in Tables 9, 9, 9.
Adv KdV SWE
Modes 10 12 12
Width 24 32 48
Layers 5 5 5
Table 7: FNO Parameters
Adv KdV SWE
Width 8 12 32
Bottleneck Dim 32 48 128
Layers 6 6 6
Table 8: Unet Parameters
Adv KdV SWE
Width 32 32 64
Kernel local global global
Cond. FiLM FiLM FiLM
Table 9: HNF Parameters
Computational Resources
All experiments were run on a single NVIDIA RTX 6000 Ada GPU; 1D experiments were usually run within 10 minutes, while 2D experiments were usually run within an hour. The largest dataset is around 20 GB, for the 2D Shallow-Water equations.